# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ 4 Experiments
We conduct our experiments in general chat and code generation. Our main setting is a single-turn general chat task, a common research testbed for offline RL fine-tuning in LLM alignment, and we add a code generation task where the model is asked to generate code that passes unit tests given a description. We focus on comparing QRPO with policy fitting methods offline, and choose DPO, REBEL, and SimPO as our baselines. Each captures a fundamental means of fitting the solution of the RL fine-tuning objective: DPO for relative preferences, REBEL for the relative reward difference, and QRPO for the absolute reward. We added SimPO to compare to methods that replace the canonical implicit reward with a reward derived from an inductive bias (length normalization). We select a comprehensive set of models, datasets, and distribution shift settings and sweep over the same comprehensive hyperparameter budget for each algorithm, resulting in 225 trained models for each algorithm. Additionally, we adopt a three-fold data split (train/validation/test) with three seeds to perform robust hyperparameter selection and report error bars. Our setting captures many of the experimental settings used in previous related work (Gao et al., 2024; Meng et al., 2024) and allows for broader comparisons. We present our experimental protocol in detail in Appendix L.
@@
Furthermore, by varying the KL regularization strength, we observe a tradeoff between using the off-policy versus the offline regime. For a relatively large regularization ($\beta=0.1$), the off-policy regime outperforms the offline regime with a relatively low number of reference rewards and continues to surpass it with more reference rewards. This aligns well with modern post-training pipelines (Grattafiori et al., 2024; Lambert et al., 2024) which use a relatively large regularization to train on large datasets targeting multiple tasks and capabilities, and perform multiple iterations of policy fitting on newly sampled data from the latest model. The Llama 3 models (Grattafiori et al., 2024), for example, have been trained with DPO with $\beta=0.1$ for several rounds, updating the reference model and sampling from it at each round; QRPO would be easily integrated into this off-policy pipeline to provide its top performance. Also note that in this case, all the previously generated reference completions can be reused as training samples to generate a growing training dataset. Yet, for a low regularization ($\beta=0.003$), the offline regime with high-quality offline data and minimal KL constraints yields the highest performance overall across $\beta$s, and the off-policy regime requires a large number of reference rewards to reach comparable performance. Although this setting combined with offline data would not provide stable results for large post-training pipelines due to a large test-time distribution shift (see Appendix A for a comprehensive discussion), it is the setting where research papers report their best results, as it eventually provides the best performance on narrow tasks, leveraging high variance combined with hyperparameter optimization. Most of the best models for all the algorithms we train also result from this low regularization offline setting, which at the same time allows QRPO to provide strong results with very few reference rewards, but we minimize the impact from a selection bias by adopting a robust model selection protocol.
Table 3: QRPO compared to baselines. We report original (Orig.) and length-controlled (LC) rewards (defined in App. L) on a held-out test split with means and standard deviations over three seeds, and the AlpacaEval 2 original and length-controlled win rates (WR and LC) with their standard errors. Higher is better. We bold the highest numbers with intersecting means $\pm$ one standard deviation after rounding. The best model for each algorithm is selected according to a validation set from training with 9 hyperparameter combinations (learning rate and $\beta$) across 6 distribution shift settings, including the option to first SFT on the chosen answer (Initial vs. SFT-chosen) and then the option to fine-tune on set dataset completions (offline) or generate new completions from the initial checkpoint (off-policy) and annotate them (using the best vs. the worst among 6 completions, or by ranking a random pair of samples). We rank the preference pairs used by DPO and SimPO using the same reward model (ArmoRM) used by REBEL and QRPO, ensuring an equal potential signal source for all methods.
@@
Llama 8B Tülu 3 SFT Initial ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1518}$$\pm$ 0.0011 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1519}$$\pm$ 0.0010 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}18.1}$$\pm$ 0.4 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}10.8}$$\pm$ 1.1 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1294}$$\pm$ 0.0006 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1293}$$\pm$ 0.0005 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}18.1}$$\pm$ 0.4 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}10.8}$$\pm$ 1.1
@@
Mistral 7B Instruct v0.2 Initial ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1598}$$\pm$ 0.0008 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1598}$$\pm$ 0.0006 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}27.6}$$\pm$ 0.4 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}21.6}$$\pm$ 1.5 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1348}$$\pm$ 0.0009 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1349}$$\pm$ 0.0009 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}27.2}$$\pm$ 0.4 ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}21.4}$$\pm$ 1.4
@@
Table 3 shows that QRPO achieves notably strong results in optimizing the offline RL fine-tuning objective for general conversational tasks, translating effectively to downstream evaluations on AlpacaEval. QRPO models consistently rank among the top-performing models based on length-controlled metrics, suggesting that the approach also mitigates the well-known length bias issue, which we analyze in the next section. Notably, on the Magpie-Air dataset—which contains completions from Llama-8B-Instruct, which is stronger than our initial models—we only use a single reference reward in QRPO to estimate the quantile reward, corroborating our results from the scaling analysis (section 4). Indeed, most of the algorithms trained on Magpie-Air obtained their best performance in the offline distribution shift setting, leveraging the quality of the dataset. For the UltraFeedback dataset, which is significantly lower quality than Magpie-Air and the initial models, we still only used 3 reference rewards for QRPO, and most of the algorithms obtained their best performance in the off-policy distribution shift setting, replacing the completions from the dataset with the reference model completions. Appendix M shows the hyperparameters and distribution shift settings for which each algorithm obtained its best model.
@@
Length bias (or correlation) is a spurious correlation between the output length and the performance of models resulting from RLHF and DAA training (Dubois et al., 2024; Singhal et al., 2024; Park et al., 2024). It’s undesired as it increases the verbosity and inference cost of the models without necessarily improving their unbiased performance. It is inherited from training on (underspecified) preference labels. However, the reward model ArmoRM (Wang et al., 2024) we use in our experiments has been explicitly trained to counteract this length bias in its scalar value and we show that QRPO and REBEL trained with this scalar value exhibit less length bias than DPO and SimPO which convert it to preferences, potentially inheriting the length bias back. Similarly to Park et al. (2024), we study the completion length versus the implicit rewards. As we have generated reference completions for each prompt to train QRPO, we are in the position to leverage a better signal for length bias than previous work. We take a per-prompt standardized completion length, which for each prompt and completion generated by the trained model, is the length of the completion standardized by the distribution of reference completions in that prompt (that we generated in the pre-computation phase). Hence, unlike prior work, our $x$-axis captures the length increase for each prompt independently. For implicit rewards, in the same manner as Park et al. (2024), for each model (policy), we consider the reward parameterized by this policy according to the RL fine-tuning objective (Eq. 3 without $Z$, colloquially the DPO reward). This is the reward for which the trained policy is optimal. Hence, if there is strong length bias or correlation in the reward, we can say that this is also a feature of the learned policy, as it is optimal for that reward. Results are shown in Figure 4 and Appendix M. We observe the typical positive slope for DPO reported by Park et al. (2024), and surprisingly, we observe the same slope for SimPO. Although SimPO decreases length overall (the cloud is shifted to the left compared to other models), the length bias trend is still present. On the contrary, we observe that QRPO and REBEL exhibit much less length bias. We conjecture that this improvement is due to effective training with the robust ArmoRM reward signal, which for SimPO and DPO is lost in the binary preferences.
@@ 5 Related work
A large body of research on policy fitting methods in RL fine-tuning has focused on developing preference-based algorithms to compete directly with DPO (Rafailov et al., 2023) by better learning from preferences. The list includes IPO (Azar et al., 2024), ORPO (Hong et al., 2024), CPO (Xu et al., 2024a), and SimPO (Meng et al., 2024). However, many of these methods deviate from optimizing the RL fine-tuning objective and attempt to better optimize human feedback for conversational abilities. With comprehensive results, SimPO positions itself as the most competitive method and shows that a well-tuned DPO is still superior to most other methods. Hence, we choose to compare to DPO to capture most of these methods, to SimPO to capture one competitive way to include inductive bias in the objective, and to REBEL to illustrate fundamentally different ways of optimizing the RL fine-tuning objective: preferences vs. reward differences.
@@
DRO (Pierre Harvey Richemond et al., 2024) is a recent approach that formulates the same regression objective presented in Eq. 6, but proceeds to learn the partition function jointly with the policy to optimize the objective. Although the approach is close to our motivation and derivation, DRO no longer benefits from the simplicity and stability of policy fitting methods, as it learns a policy and a value function through joint optimization. In this work, we evaluate simple policy fitting methods and thus do not compare against DRO (which also does not have an open implementation). KTO (Ethayarajh et al., 2024) simply drops the partition function in the KL-regularized optimal reward (Eq. 3) and optimizes for human utility based on prospect theory (Kahneman & Tversky, 2013) instead of preferences, making the partition function irrelevant and the objective optimizable with pointwise binary human feedback. KTO could be considered as a baseline using a pointwise signal; but, since it excessively diverges from the RL fine-tuning objective, and Meng et al. (2024) show that empirically, using similar benchmarks to us, KTO is less performant than DPO and SimPO, we opted not to include KTO as a baseline for comparison.
@@
The cumulative density function of the rewards under the reference policy has been used in both the vBoN (Amini et al., 2025) and BOND (Sessa et al., 2025) algorithms to derive the distribution of the best-of-N (BoN) policy and identify the BoN policy as induced by optimizing the $\log$-quantile rewards in the RL fine-tuning objective. However, both derivations focus specifically on the event defining the BoN probability and do not stem from the solution to the KL-regularized RL objective, which in these works remains intractable due to the partition function. QRPO instead uses the quantile reward as a universal solution to derive the partition function and the expression of the induced policy. In fact, we have the BoN policy as a special case of our framework using the $\log$-quantile transformation. Another difference with our work is that we opt for a regression loss rather than a distribution matching loss. Gulcehre et al. (2023, ReST) also experiment with an SFT loss on the best off-policy data according to a quantile reward threshold. We believe that comparing QRPO, BOND, vBoN, and ReST can be made in the scope of the more targeted open question of studying the BoN distillation problem, but is out of the scope of the claims made in our work.
@@
There is a growing body of research attempting to make policy improvement methods work off-policy, such as Roux et al. (2025, TOPR), which uses truncated importance sampling ratios. This is a promising direction, confirming that QRPO tackles an important gap in the literature to optimize pointwise absolute rewards with data from other distributions.

### v2

@@ 4 Experiments
We conduct our experiments in general chat and code generation. Our main setting is a single-turn general chat task, a common research testbed for offline RL fine-tuning in LLM alignment, and we add a code generation task where the model is asked to generate code that passes unit tests given a description. We focus on comparing QRPO with policy fitting methods offline, and choose DPO, REBEL, and SimPO as our baselines. Each captures a fundamental means of fitting the solution of the RL fine-tuning objective: DPO for relative preferences, REBEL for the relative reward difference, and QRPO for the absolute reward. We added SimPO to compare to methods that replace the canonical implicit reward with a reward derived from an inductive bias (length normalization). We select a comprehensive set of models, datasets, and distribution shift settings and sweep over the same comprehensive hyperparameter budget for each algorithm, resulting in 225 trained models for each algorithm. Additionally, we adopt a three-fold data split (train/validation/test) with three seeds to perform robust hyperparameter selection and report error bars. Our setting captures many of the experimental settings used in previous related work (gao_rebel_2024; meng2024simpo) and allows for broader comparisons. We present our experimental protocol in detail in Appendix M.
@@
[…]
Table 3: QRPO compared to baselines. We report original (Orig.) and length-controlled (LC) rewards (defined in App. M) on a held-out test split with means and standard deviations over three seeds, and the AlpacaEval 2 original and length-controlled win rates (WR and LC) with their standard errors. Higher is better. We bold the highest numbers with intersecting means $\pm$ one standard deviation after rounding. The best model for each algorithm is selected according to a validation set from training with 9 hyperparameter combinations (learning rate and $\beta$) across 6 distribution shift settings, including the option to first SFT on the chosen answer (Initial vs. SFT-chosen) and then the option to fine-tune on set dataset completions (offline) or generate new completions from the initial checkpoint (off-policy) and annotate them (using the best vs. the worst among 6 completions, or by ranking a random pair of samples). We rank the preference pairs used by DPO and SimPO using the same reward model (ArmoRM) used by REBEL and QRPO, ensuring an equal potential signal source for all methods.
@@
Llama 8B
Tülu 3
SFT
Initial $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1518$} $\pm$ 0.0011 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1519$} $\pm$ 0.0010 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}18.1$} $\pm$ 0.4 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}10.8$} $\pm$ 1.1 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1294$} $\pm$ 0.0006 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1293$} $\pm$ 0.0005 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}18.1$} $\pm$ 0.4 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}10.8$} $\pm$ 1.1
@@
Mistral 7B
Instruct
v0.2
Initial $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1598$} $\pm$ 0.0008 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1598$} $\pm$ 0.0006 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}27.6$} $\pm$ 0.4 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}21.6$} $\pm$ 1.5 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1348$} $\pm$ 0.0009 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1349$} $\pm$ 0.0009 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}27.2$} $\pm$ 0.4 $\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}21.4$} $\pm$ 1.4
@@
Table 3 shows that QRPO achieves notably strong results in optimizing the offline RL fine-tuning objective for general conversational tasks, translating effectively to downstream evaluations on AlpacaEval. QRPO models consistently rank among the top-performing models based on length-controlled metrics, suggesting that the approach also mitigates the well-known length bias issue, which we analyze in the next section. Notably, on the Magpie-Air dataset—which contains completions from Llama-8B-Instruct, which is stronger than our initial models—we only use a single reference reward in QRPO to estimate the quantile reward, corroborating our results from the scaling analysis (section 4). Indeed, most of the algorithms trained on Magpie-Air obtained their best performance in the offline distribution shift setting, leveraging the quality of the dataset. For the UltraFeedback dataset, which is significantly lower quality than Magpie-Air and the initial models, we still only used 3 reference rewards for QRPO, and most of the algorithms obtained their best performance in the off-policy distribution shift setting, replacing the completions from the dataset with the reference model completions. Appendix N shows the hyperparameters and distribution shift settings for which each algorithm obtained its best model.
@@
Length bias (or correlation) is a spurious correlation between the output length and the performance of models resulting from RLHF and DAA training (dubois2024length; singhallong; park2024disentangling). It’s undesired as it increases the verbosity and inference cost of the models without necessarily improving their unbiased performance. It is inherited from training on (underspecified) preference labels. However, the reward model ArmoRM (wang_interpretable_2024) we use in our experiments has been explicitly trained to counteract this length bias in its scalar value and we show that QRPO and REBEL trained with this scalar value exhibit less length bias than DPO and SimPO which convert it to preferences, potentially inheriting the length bias back. Similarly to park2024disentangling, we study the completion length versus the implicit rewards. As we have generated reference completions for each prompt to train QRPO, we are in the position to leverage a better signal for length bias than previous work. We take a per-prompt standardized completion length, which for each prompt and completion generated by the trained model, is the length of the completion standardized by the distribution of reference completions in that prompt (that we generated in the pre-computation phase). Hence, unlike prior work, our $x$-axis captures the length increase for each prompt independently. For implicit rewards, in the same manner as park2024disentangling, for each model (policy), we consider the reward parameterized by this policy according to the RL fine-tuning objective (Eq. 3 without $Z$, colloquially the DPO reward). This is the reward for which the trained policy is optimal. Hence, if there is strong length bias or correlation in the reward, we can say that this is also a feature of the learned policy, as it is optimal for that reward. Results are shown in Figure 4 and Appendix N. We observe the typical positive slope for DPO reported by park2024disentangling, and surprisingly, we observe the same slope for SimPO. Although SimPO decreases length overall (the cloud is shifted to the left compared to other models), the length bias trend is still present. On the contrary, we observe that QRPO and REBEL exhibit much less length bias. We conjecture that this improvement is due to effective training with the robust ArmoRM reward signal, which for SimPO and DPO is lost in the binary preferences.
@@ 5 Related work
A large body of research on policy fitting methods in RL fine-tuning has focused on developing preference-based algorithms to compete directly with DPO (rafailov2023direct) by better learning from preferences. The list includes IPO (azar2023ipo), ORPO (hong2024orpo), CPO (xu2024contrastive), and SimPO (meng2024simpo). However, many of these methods deviate from optimizing the RL fine-tuning objective and attempt to better optimize human feedback for conversational abilities. With comprehensive results, SimPO positions itself as the most competitive method and shows that a well-tuned DPO is still superior to most other methods. Hence, we choose to compare to DPO to capture most of these methods, to SimPO to capture one competitive way to include inductive bias in the objective, and to REBEL to illustrate fundamentally different ways of optimizing the RL fine-tuning objective: preferences vs. reward differences.
@@
DRO (pierre_harvey_richemond_offline_2024) formulates the same regression objective in Eq. 6, but proceeds to learn the partition function jointly with the policy. Although the approach is close to our motivation and derivation, DRO no longer benefits from the simplicity and stability of policy fitting methods, as it learns a policy and a value function through joint optimization. In this work, we evaluate simple policy fitting methods and thus do not compare against DRO (which also does not have an open implementation). More recently, the same objective has also been used in SPO (cohen2025soft), AGRO (tang2025rl), and TBA (bartoldson2025trajectory) which like DRO get both the policy and the partition function from Eq. 5 but instead of learning the partition function they propose different ways of estimating it. These works mostly focus on semi-online training and do not demonstrate competitive performance in the offline regime. Alternatively, KTO (ethayarajh_kto_2024) drops the partition function in the KL-regularized optimal reward (Eq. 3) and optimizes for human utility based on prospect theory (kahneman2013prospect) instead of preferences, making the partition function irrelevant and the objective optimizable with pointwise binary human feedback. KTO could be considered as a baseline using a pointwise signal; but, since it excessively diverges from the RL fine-tuning objective, and meng2024simpo show that empirically, using similar benchmarks to us, we do not KTO consider as a baseline.
@@
The cumulative density function of the rewards under the reference policy has been used in both the vBoN (amini2025variational) and BOND (sessa2025bond) algorithms to derive the distribution of the best-of-N (BoN) policy and identify the BoN policy as induced by optimizing the $\log$-quantile rewards in the RL fine-tuning objective. However, both derivations focus specifically on the event defining the BoN probability and do not stem from the solution to the KL-regularized RL objective, which in these works remains intractable due to the partition function. QRPO instead uses the quantile reward as a universal solution to derive the partition function and the expression of the induced policy. In fact, we have the BoN policy as a special case of our framework using the $\log$-quantile transformation. Another difference with our work is that we opt for a regression loss rather than a distribution matching loss. gulcehre2023reinforced also experiment with an SFT loss on the best off-policy data according to a quantile reward threshold. We believe that comparing QRPO, BOND, vBoN, and ReST can be made in the scope of the more targeted open question of studying the BoN distillation problem, but is out of the scope of the claims made in our work.
@@
There is a growing body of research attempting to make policy improvement methods work off-policy, such as TOPR (roux2025tapered), which uses truncated importance sampling ratios, and ReMix (liang2025squeeze), which combines regularization techniques for data reuse. This is a promising direction, confirming that QRPO tackles an important gap in the literature to optimize pointwise absolute rewards with data from other distributions.